OpenStudy (anonymous):

Find the derivatives. p(x)=(3x^2-1)(7+2x^3)

OpenStudy (cggurumanjunath):

use product rule @malibugranprix2000 !

OpenStudy (anonymous):

I see that the product rule is used, but I am not sure how to figure it out still?

OpenStudy (cggurumanjunath):

(uv)'=uv'+vu'

OpenStudy (cggurumanjunath):

u-(3x^2-1) v-(7+2x^3)

OpenStudy (cggurumanjunath):

(3x^2-1)(7+2x^3)'+(7+2x^3)(3x^2-1)'

OpenStudy (cggurumanjunath):

'-derivative

OpenStudy (cggurumanjunath):

(7+2x^3)'=(0+2(3x^2))=6x^2

OpenStudy (cggurumanjunath):

(3x^2-1)'=(3*2x)

OpenStudy (cggurumanjunath):

is it clear @malibugranprix2000 ?

OpenStudy (cggurumanjunath):

can u proceed from here ?

OpenStudy (anonymous):

@CGGURUMANJUNATH, yes I will work on it.

OpenStudy (cggurumanjunath):

very good.

OpenStudy (cggurumanjunath):

if you are not clear at any step kindly ask me . @malibugranprix2000

OpenStudy (anonymous):

@CGGURUMANJUNATH thanks for the help so far :)

OpenStudy (cggurumanjunath):

:)